A firm has 33 units of a certain product on hand. Forecasts for the first two planning periods are 20 units each. A production quantity of 20 units is planned to be available in period 2. What is the projected on-hand inventory at the end of period 2?
Answer #1

Projected on hand inventory at the end of period 1 = Beginning Inventory - Forecast + scheduled receipts in period 1

Projected on hand inventory at the end of period 1 = 33 - 20 + 0 = 13

Projected on hand inventory at the end of period 2 = Ending Inventory of period 1 - Forecast + scheduled receipts in period 2

Projected on hand inventory at the end of period 2 = 13 - 20 + 20 = 13 units